Complete the identity below:
(csc (x+1)(csc x-1)
__________________ = ?
cot^2 x

Found 2 solutions by stanbon, gsmani_iyer:
Answer by stanbon(75887)

[(csc x+1)(csc x-1)] / cot^2 x = ?
-------------------------
(csc^2(x)-1)/cot^2(x) = x
-----
Since csc^2(x) -1 = cot^2(x),
-----
cot^2(x)/cos^2(x) = 1
====================
Cheers,
Stan H.
---------------

Answer by gsmani_iyer(201)


By simplification, the numerator equals csc^2 x - 1 which is equals to cot^2 x.
Therefore, (cot^2 x) upon (cot^2 x) = 1
I hope this is clear enough. Feel free to contact for further doubts.